# UrbanVLP: Multi-Granularity Vision-Language Pretraining for Urban Region Profiling

### v2

@@ Appendix E More Details for Modality Representation
For utilizing ViT as the visual encoder to process urban visual data, in particular, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT. Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers. Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection. We also prepend a learnable embedding to the patch embedding sequence (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$). Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information. The process is as follows:
@@
where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$ and $\mathbf{E}_{pos}\in\mathbb{R}^{(N+1)\times d}$ are learnable parameters. $\mathbf{z}_{0}$ represents the result of the linear projection, which is subsequently fed into the ViT.
Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation [68]. Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are after every block:
@@
Appendix F Segmentation Ratio and Prompt Examples
As illustrated in Figure 15(a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityScape}$ dataset. The segmentation ratio is decoupled by pretrained segmentation models and is used in both prompt construction and PerceptionScore computation. We group the segmentation categories into 13 categories following [17]: Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
The prompt template used to generate street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner. The coordinate of the street-view image is [Longitude, Latitude]. The segmentation ratio of the street-view image is [Segment Information]. Figure 15(c) provides illustrations of some examples of our prompts used for description generation.
@@ Appendix F Segmentation Ratio and Prompt Examples
Appendix G The Choice of Image-to-Text LMM
In our work, we leverage ShareGPT4V [11] for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table LABEL:table:lmms. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in [74]. Additionally, we note that there is study [86] argue that InstructBLIP [14] exhibits state-of-the-art (SOTA) performance in image-to-text generation. However, in our experiment, we found that InstructBLIP is unsuitable for handling geospatial coordinates as prompt content. Incorporating such coordinates significantly degrades the quality of the generated text, leading to poor content output. Due to these limitations, we have excluded it from our comparative analysis.
Table 3: Perception score of representative LMMs.
Models Average Generated-Description Scores
LLaVA-v1.5 [48] 0.674
mPLUG-Owl [77] 0.662
LLaMA-Adapter V2 [20] 0.698
MiniGPTv2 [10] 0.693
ShareGPT4V [11] 0.714
Appendix H More Experimental Analysis
H.1 More Analysis for Text Calibration
In our paper, we selected the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning. In order to ascertain the efficacy of ShareGPT4V, we computed the distribution of PerceptionScore for descriptions generated by ShareGPT4V, as depicted in Figure 14. Notably, only less than 0.02% of generated texts scored below 0.5, thereby evidencing the high quality of the descriptions produced.
Table 4: UrbanVLP performance utilizing mPLUG-Owl with and without Text Calibration. mPO is an abbreviation for mPLUG-Owl, and TC stands for Text Calibration.
@@ H.1 More Analysis for Text Calibration
As a result, the experimental results comparing with and without calibration do not adequately reflect the effectiveness of our calibration method. To further demonstrate the effectiveness of our method, we select mPLUG-Owl [77], one of the representative LMMs compared in our paper, with the worst performance. After calibration, 15.3% of the text with perceptionscores below 0.5 was filtered out. The results validate the effectiveness of our calibration method, and indicate that more relevant and noise-free textual information could align better with image features, leading to a more coherent and meaningful visual representation.
H.2 Analysis about Time Complexity
@@
Modeling and research in urban areas can provide valuable references and assistance to policymakers, builders, and citizens. Therefore, practical implementation should be considered. We conduct time complexity analysis using FPS (Frames Per Second), a common indicator for inference speed in computer vision area. We report FPS for satellite image in Figure 16(b). In fact, introducing street view features for better and finer-grained effects will inevitably sacrifice a portion of the time efficiency. However, compared to [41]’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance. For making model decisions during deployment, to pursue the trade-off between performance and speed, UrbanVLP is the optimal choice.
Figure 16: (a) Hyperparameter for loss combination. (b) FPS comparison of different mdoels. (c) Expetiments for maximum of the number of street-view images.
H.3 Analysis about Hyperparameters
The optimal hyperparameters for Equation 7 are actually determined by grid search. In Figure 16(a), we present some results of different combinations of hyperparameter values . It can be seen that the settings we chose in the paper ($\alpha$=0.5, $\beta$=0.5) yield the best experimental results.
We also study the maximum value of street-view images during training, which is set as 25 in our paper. As is shown in Figure 16(c), as the number of street-view images increases, the R² value continuously rises, reaching its peak at 25. Therefore, our setting is the optimal choice.
Appendix I Practicality
@@
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban indicator framework using the Mapbox platform [56]. This system renders a visual representation of urban areas through satellite imagery, complemented by street-level photographs for an interactive exploration experience. Figure 17 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
Each urban region captured by a satellite image, encompassing a 1km $\times$ 1km area, is marked by a blue dot, which is encircled by yellow points that denote the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.
Furthermore, a street-view panel allows users to peruse individual street-view images alongside pertinent captions. The system also facilitates the discovery of popular POIs within any given area, enriching the understanding of its functional characteristics. Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators.
Appendix J Limitations and Social Impact
Due to the easier accessibility of the dataset, our work primarily focuses on first-tier and domestic cities, which introduces a certain bias and can be regarded as a limitation of our study. Moreover, the model’s performance is significantly influenced by image resolution; higher quality image data is required to achieve optimal performance. Additionally, the introduction of new modalities inevitably leads to a decrease in inference efficiency, which should be a consideration for practical deployment. Moreover, the inability to automatically correct low-quality text highlights an area for future improvement.
Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding urban planning and decision-making processes. However, it is crucial to note that satellite and street-view images might inadvertently contain private information. The street-view images we collected have been blurred by the Baidu Map to obscure relevant information, such as license plates, thereby significantly reducing associated risks. Nevertheless, we strongly advise users to perform a comprehensive analysis of the data prior to deploying our model in practical contexts.

### v3

@@ G    More Details for Modality Representation
For utilizing ViT as the visual encoder to process urban visual data, we reshape the input image $I\in\mathbb{R}^{H\times W\times 3}$ into a sequence of flattened 2D patches ${I}_{P}\in\mathbb{R}^{N\times({P^{2}}\cdot{C})}$, where C represents the number of channels, $(P,P)$ specifies the resolution of each image patch, and $N=HW/P^{2}$ denotes the number of patches, corresponding to the input sequence length for ViT. Notebly, the ViT utilizes a consistent latent vector size (denoted as $d$) across all its layers. Thus, we can obtain the outputs as the patch embeddings through a trainable linear projection. We also prepend a learnable embedding to the patch embedding sequence to serve as the overall representation (i.e., $\mathbf{z}_{0}^{0}={I}_{cls}$). Besides, we add learnable position embedding $\mathbf{E}_{pos}$ to retain positional information. The process is as follows:
@@
where $\mathbf{W}\in\mathbb{R}^{(P^{2}\cdot C)\times d}$
Consistent with the standard Transformer encoder, ViT includes alternating layers of multi-head self-attention $\operatorname{MSA(\cdot)}$ operation (Vaswani et al. 2017). Furthermore, layer normalization (LN) is adopted before every block, whereas residual connections are applied after every block:
@@
H    Segmentation Ratio and Prompt Examples
As illustrated in Figure 15 (a) and (b), we demonstrate examples of the segmentation ratio of street-view images in our $\mathtt{CityView}$ dataset. The segmentation ratio is decoupled by pretrained segmentation models and is used in both prompt construction and PerceptionScore computation. We group the segmentation categories into 13 categories following (Fan et al. 2023b): Person; Bike; Heavy Vehicle; Light Vehicle; Facade; Window & Opening; Road; Sidewalk; Street Furniture; Greenery - Tree; Greenery - Grass & Shrubs; Sky; Nature.
The prompt template for generating street-view descriptions is formulated as follows: Analyze the street-view panoramic image in [City] in a comprehensive and detailed manner. The coordinate of the street-view image is [Longitude, Latitude]. The segmentation ratio of the street-view image is [Segment Information]. Figure 15 (c) provides illustrations of some examples of our prompts for description generation.
@@ H    Segmentation Ratio and Prompt Examples
I    The Choice of Image-to-Text LMM
In our work, we leverage ShareGPT4V (Chen et al. 2023b) for text description generation. Here we conduct experiments on current open-sourced LMM models, and compute the PerceptionScore of them. The results are presented in Table 3. As we can see, ShareGPT4V achieves the best performance in terms of PerceptionScore, followed by LLaMA-Adapter V2, which is used as the Image-to-Text models in (Yan et al. 2024). Additionally, we note that there is study (Zhong et al. 2024) argue that InstructBLIP (Dai et al. 2023) exhibits state-of-the-art (SOTA) performance in image-to-text generation. However, in our experiment, we found that InstructBLIP is unsuitable for handling geospatial coordinates as prompt content. Incorporating such coordinates significantly degrades the quality of the generated text, leading to poor content output. Due to these limitations, we have excluded it from our comparative analysis.
Table 3: PerceptionScore of representative LMMs. GD is short for Generated-Description.
Models Average GD-Scores
LLaVA-v1.5 (Liu et al. 2023a) 0.674
mPLUG-Owl (Ye et al. 2023) 0.662
LLaMA-Adapter V2 (Gao et al. 2023) 0.698
MiniGPTv2 (Chen et al. 2023a) 0.693
ShareGPT4V (Chen et al. 2023b) 0.714
J    More Experimental Analysis
More Analysis for Text Calibration
In our paper, we select the state-of-the-art LMM ShareGPT4V, which demonstrates performance comparable to GPT4V in image captioning. Here, we compute the distribution of PerceptionScore for descriptions generated by ShareGPT4V in Figure 16. Moreover, in our proposed dataset, we filter out the texts with PerceptionScore lower than 0.6. To validate the effectiveness of this Text Calibration process, we report the performance of using raw generated descriptions (i.e., UrbanVLP w/o Text Calibration) for comparison.
Table 4 demonstrates that UrbanVLP consistently surpasses this variant across various indicators. This outcome confirms the effectiveness of our Text Calibration method, suggesting that aligning more relevant and noise-free textual information with image features enhances coherence and meaning in visual representation.
Figure 16: PerceptionScore distribution on $\mathtt{CityView}$ dataset for 4 representative cities.
Table 4: UrbanVLP performance with and without Text Calibration. TC stands for Text Calibration.
@@ I    The Choice of Image-to-Text LMM
Analysis about Time Complexity
Researches in urban areas can provide valuable references and assistance to policymakers, builders, and citizens. Therefore, the practicality of the model should be taken into consideration.
@@
We conduct time complexity analysis using FPS (Frames Per Second), a common indicator for inference speed in computer vision area. The result is reported in Figure 17 (a). In fact, compared to UrbanCLIP, UrbanVLP’s introduction of street-view features for improved and finer-grained effects will inevitably sacrifice a portion of the time efficiency. However, compared to StructuralUrban’s graph structure inference, our linear probing setting ensures the minimum time overhead, also with better performance. For making model decisions during deployment, to pursue the trade-off between performance and speed, UrbanVLP is the optimal choice. We also present a practical deployment example of UrbanVLP in Appendix K.
Figure 17: (a) FPS comparison of different models. (b) Comparison of the quality of generated texts of UrbanVLP and UrbanCLIP.
Discussion on the Enhancements in Text Generation Quality by UrbanVLP
As we discussed in the Introduction, the text generation process of UrbanCLIP raises several critical concerns: i) Hallucination and ii) Homogenization. To assess the comparative improvement of UrbanVLP in addressing these issues, we calculate three metrics for the generated text in both UrbanCLIP and UrbanVLP: (1) the average text length, (2) the mean text similarity, and (3) the average PerceptionScore.
First of all, the average length of the text description in UrbanCLIP is quantified at 18.1 words. This relatively short length may lead to insufficient information being conveyed. In contrast, the average length of the text description in UrbanVLP is 144.1 words, nearly eight times longer, providing ample space to describe the content of the images in greater detail, thereby covering more nuanced information. Furthermore, we calculate the mean text similarity of all the generated texts of UrbanCLIP and UrbanVLP. As depicted in Figure 17 (b) on the left side, the mean text similarity of UrbanCLIP texts is 0.28, while that of UrbanVLP is 0.19. This indicates that our Automatic Text Generation and Calibration mechanism significantly mitigates the issue of homogenization, resulting in more diversified text descriptions for images with different content. Additionally, on the right side of Figure 17 (b), we evaluate the quality of generated descriptions through our proposed PerceptionScore. As we can see, the average scores of UrbanCLIP texts and UrbanVLP texts are 0.675 and 0.714, respectively. This indicates that UrbanVLP effectively mitigates the issue of inconsistency (hallucination) between generated text and visual content.
Ablation study for fusion method
To verify the effectiveness of our addition fusion method of satellite image features, street-view image features and location features, we experiment with different types fusion method, including (1) Concatenation: we directly concat the three kinds of features; (2) MLP: we feed the three features into a multi-layer perceptron (MLP) for fusion, resulting in a single output. As indicated in Table 5, the differences among several distinct fusion methods are relatively minor, which may be attributed to the overall complexity of the model, resulting in minimal impact from the various fusion techniques. Consequently, we opt for the most straightforward and high-performing approach, the addition method, for implementation.
Table 5: Ablation study on $\mathtt{CityView}$-$\mathtt{Beijing}$ dataset. The best results are in bold.
Model Carbon Population GDP
$\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE $\text{R}^{2}$ RMSE MAE
Concat 0.759 0.487 0.389 0.705 0.535 0.429 0.526 0.693 0.424
MLP 0.761 0.483 0.374 0.710 0.527 0.415 0.532 0.688 0.418
Addition (Ours) 0.769 0.477 0.369 0.714 0.523 0.411 0.537 0.684 0.416
More Visualization of Region Representations
Figure 18 provides a visualization with a population indicator, compared to UrbanCLIP. We use the dot color to indicate the population (log) in the corresponding regions of Beijing. As shown by the red dashed boxes, the regions modeled by UrbanVLP exhibit more significant clustering characteristics and greater distinction in high-dimensional space. Additionally, the lower parts of the two main band-shaped branches are also more distinctly separated. Points that are spatially close also have similar population values.
Figure 18: PCA visualization of urban imagery.
K    Practicality
@@
To demonstrate the practical applications of our model in real-life scenarios, we have developed a system bearing the same name UrbanVLP. The UrbanVLP system is a novel web-based system developed to integrate LMMs with a multi-scale urban socioeconomic indicator framework using the Mapbox platform (Mapbox 2024). Figure 19 showcases the user-friendly interface of the map, highlighting functionalities that facilitate zooming, searching for locations, and navigating through various sectors.
Each urban region, captured by a satellite image and covering a 1km $\times$ 1km area, is marked by a blue dot. Yellow points surround the blue dot, indicating the locations of street-view images accessible within that region. Engaging with these points allows users to uncover a suite of urban metrics, such as carbon emissions, population, GDP, night light, house price, and POI. The system’s visual components are augmented by a captioning feature that provides clear and informative textual descriptions, offering insights into the spatial characteristics of the designated area.
Furthermore, a street-view panel allows users to peruse individual street-view images alongside pertinent captions. The system also facilitates the discovery of popular POIs within any given area, enriching the understanding of its functional characteristics. Ultimately, UrbanVLP is designed to distill complex urban data into a visual and user-friendly format, thereby offering a holistic and detailed view of urban landscapes and their key indicators. We have also prepared a demonstration video in the supplementary material to showcase the utility of our system.
L    Limitations and Social Impact
Due to the easier accessibility of the data, our work primarily focuses on Chinese first-tier cities, which introduces a certain bias and can be regarded as a limitation of our study. Moreover, the model’s performance is significantly influenced by image resolution. Higher quality image data is required to achieve optimal performance. Additionally, the inability to automatically correct low-quality text highlights an area for future improvement.
Our work can be deployed in practical applications, as demonstrated by the example provided in the paper, thereby aiding in urban planning and decision-making processes. However, it is crucial to note that satellite and street-view images might inadvertently contain private information. The street-view images we collected have been blurred to obscure relevant information, such as license plates, thereby significantly reducing associated risks. Nevertheless, we strongly advise users to perform a comprehensive analysis of the data prior to deploying our model in practical contexts.